OpenStudy (anonymous):

A sports team is building a new stadium on a rectangular lot of land. If the lot measures 6x by 10x and the sports field will be 1x by 4x, how much of the lot will be left over to build bleachers on? 56x² 64x² 30x² 60x² – 4x²

OpenStudy (rulnick):

60x^2-4x^2 = 56 x^2

OpenStudy (anonymous):

6x by 10x = 60 x^2 1x by 4x = 4x^2 -> Left = 56x^2